Assume that we want to come concert De ce este de collected as well as the team starter Question to 1 CA choice c is incorrect. could someone help me solve this and explain how?
Answer #1

Solution:

Given that

confidence level = c = 90% = 0.90

sample size = n = 53

Since the population standard deviation is unknown , we use t distribution.

(when it is known , we use z distribution)

c = 0.90

\therefore\alpha = 1- c = 1- 0.90 = 0.10

\therefore  \alpha/2 = 0.10 \slash 2 = 0.05

Also, d.f = n - 1 = 53 - 1 = 52

\therefore  ta/2.d.f-  =  ta/2.n-1  =  t0.05,52 = 1.674689 = 1.67

(use t table or t value calculator)

Answer is

A. ta/2.d.f- = 1.67